OpenStudy (anonymous):

Find the set of values of y for which y^2(y^2-5)>36

OpenStudy (anonymous):

try distributing (multiplying) the y^2 at the front across the parenthesis... In other words, simplify the expression first.

OpenStudy (anonymous):

Or, never mind... don't do that

OpenStudy (anonymous):

ok, let me start over :) Simplify the left side of the equation by multiplying... then subtract 36 from both sides to get -36 on the left. Then you can factor the expression on the left.

OpenStudy (anonymous):

y^4-5y^2-36>0??

OpenStudy (anonymous):

yes :) Then you can factor... just think if it was y^2 - 5y - 36.... you'd factor as (y-9)(y+4) So this is the same, but you have y^4 and y^2 terms instead of y^2 and y

OpenStudy (anonymous):

I get (y^2 - 9)(y^2 + 4) > 0 So y^2 - 9 must be greater than 0 and y^2 + 4 must be greater than 0 ...or else both must be less than zero so their product multiplies to be positive.

OpenStudy (anonymous):

you need y > 3 or y < -3... then y^2 will be greater than 9 making the first parenthesis positive, and the second parenthesis will still be positive, so the inequality is still true.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

How did you factor y^4-5y^2-36>0

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

try this... it will seem like a trick, but that's ok... let some other variable by equal to y^2... let's call it "b"... so b = y^2

OpenStudy (anonymous):

then let's rewrite the expression we are trying to factor using b in place of y^2 b^2 - 5b - 36

OpenStudy (anonymous):

at that point, I thought to myself, what factors can you multiply to get 36? 6 and 6, but then the middle term would be -5b 12 and 3... same problem hey... 9 and 4 work, if the signs are correct... b^2 - 5b - 36 = (b - 9)(b + 4)

OpenStudy (anonymous):

is that making sense so far? After factoring the easier looking expression using b, then you can just put the y^2 back in, replacing the b terms in the factored expression... (b - 9)(b + 4) --->>> (y^2 - 9)(y^2 + 4)

OpenStudy (anonymous):

Sort of just one question

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

@Hollywood_chrissy what was your question?

OpenStudy (anonymous):

You said let y^2=b. But when you set up the expression you had a b^2.

OpenStudy (anonymous):

yes, that's because the original expression had a y^4 term... so b = y^2 isn't enough y^4 = (y^2)(y^2) = (b)(b) = b^2

OpenStudy (anonymous):

Ok cool

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

You said let y^2=b. But when you set up the expression you had a b^2.